In order for the mass traveling around a circular roller coaster loop to stay on the track at the top of the loop, what must the mass' minimum speed be at the top of the 10 m radius loop?
Answer #1

mg = mv^2/r

So, v^2 = gr

Vmin = sqrt(gr) = sqrt(9.8 * 10) = 9.89 m/s
Answer #2

v^2/r= g

v =sqrt(rg)

= sqrt(98.1)

= 9.9045m/s